OpenStudy (scarlettfarra2000):

Simplify the expression –4(x – 7) + 3(x + 9).

OpenStudy (anonymous):

Let's simplify step-by-step. −4(x−7)+3(x+9) Distribute: =(−4)(x)+(−4)(−7)+(3)(x)+(3)(9) =−4x+28+3x+27 Combine Like Terms: =−4x+28+3x+27 =(−4x+3x)+(28+27) =−x+55 Answer: =−x+55 Hope this helps :D

OpenStudy (scarlettfarra2000):

Thank you you have no idea how much that helps

OpenStudy (anonymous):

np :D

OpenStudy (anonymous):

if you have any questions like that then this website will show you how to do it step by step :D http://www.mathpapa.com/algebra-calculator.html